# ArtCrafter: Text-Image Aligning Style Transfer via Embedding Reframing

### v1

@@ 1 Introduction
Diffusion-based text-to-image generations [33, 29] have achieved notable progress in the realms of personalization and customization, especially in consistent synthesis such as identity protection [23, 44], object customization [15, 27] and style transfer [6, 52, 5, 8, 53, 20, 16]. In particular, text-guided style transfer focuses on fine-grained style representation that encompasses abstract concepts like texture, color, composition, and material, to create a spectrum of personalized outputs grounded in the textual semantic essence.
@@
Current methods often use pre-trained diffusion models [25, 11, 48, 10, 42, 24, 35, 43, 17] for stylization tasks, enhancing model features by adding a trainable adapter module without full retraining. In text-to-image style transfer applications, adapter-based methods shape the style and content of the output by adjusting the condition guidance scales over the input image and text prompts. However, we have identified three main issues with previous research: 1) Inadequate artistic style representation. For convenience, traditional image encoder architectures and training data primarily focus on natural images, which limits their ability to capture the intricate textures and stylistic nuances of artistic images. 2) The text-guided conditions underperform expectations. As shown in Fig. 2, the usual adapter-based [48] method fails to deliver the expected results when the text condition “Fashion shoes” is used. This discrepancy arises because the amount of information in image and text embeddings are not equivalent, yet adapter-based methods [25, 11, 48, 10, 42, 24, 35] often directly concatenate the two without addressing the imbalance and disparity, leading to image data overshadowing text prompts during the sampling process. 3) Lack of output diversity. The constrained liberation of textual guidance results in generated outputs that closely resemble the reference images, thereby limiting the diversity of the results.
@@
To tackle the aforementioned issues, we introduce ArtCrafter, a crafted embedding reframing solution based on diffusion models, designed specifically for text-guided stylization tasks. The proposed ArtCrafter consists of three components: 1) Attention-based style extraction (Sec. 3.2) leverages multi-level features to capture complicated stylistic information, ensuring more coherent and accurate stylistic encoding. It employs perceiver attention and a multi-layer architecture to capture both local and global stylistic elements. Additionally, we introduce the ArtMarket dataset, which pairs art images with text descriptions, allowing us to fine-tune an encoder initially trained on natural images. This approach retains strong generalization capabilities while being sensitive to the unique visual elements of artistic styles. 2) The text-image aligning augmentation (Sec. 3.3) module improves the alignment of image and text embeddings through crafted attentional interactions, mapping them into a shared feature space. This alignment ensures that the generated images reflect both the style of the reference image and the content of the textual conditions, thus enhancing controllability. 3) The explicit modulation component (Sec. 3.4) enhances the adaptability of conventional fusion techniques through the implementation of linear interpolation and concatenation schemes. This method facilitates the merging of original image and text embeddings with multimodal embeddings, yielding enhanced embeddings. It allows ArtCrafter to generate images that are relevant to the text prompts and exhibit diverse visual representations. Our main contributions can be summarized as follows:
@@
  • We present ArtCrafter, a lightweight adapter designed to enhance the capabilities of pre-trained diffusion models in the domain of text-guided stylized image generation. Our approach centers on attention-based style feature extraction, adeptly capturing both local and global features.
@@
  • We propose an innovative text-image aligning augmentation module that facilitates robust interaction between reference images and textual descriptions within a shared feature space, significantly amplifying the impact of text prompts on the generative process.
@@
  • The explicit modulation within ArtCrafter boosts the multimodal embeddings utilization while surpassing conventional methods in flexibility and diversity. Furthermore, ArtCrafter is compatible with additional control conditions and outperforms the state-of-the-art approaches across various experimental benchmarks.
@@
2 Related Work
@@
Attention Control in diffusion models. Following the remarkable progress made in the field of pre-training text-to-image diffusion models [29, 33, 38, 14], a series of image editing efforts [50, 2, 46, 1] have emerged. Hertz et al. [12] propose the Prompt-to-Prompt method, which achieves text-based partial image editing and generates edited images that conform to textual conditions by replacing original vocabulary and cross-attention maps. Plug-and-play [49] utilizes the spatial features and self-attention mapping of the original image to guide the diffusion model for text-guided image-to-image translation while preserving the spatial layout of the original image. Later, MasaCtrl [4] proposes a mutual self-attention control technique for coherent image editing. Consistent image editing is achieved by preserving the key and value of the self-attention layer of the source image while conditioning the model with desired text prompts. Recently, StyleID [6] proposes a style migration method without training by injecting styles in the self-attention layer and introduces query preservation, attention temperature scaling, and initial latent AdaIN techniques to minimize the impact of style injection on the original content. $\mathcal{Z}^{*}$ [8] shows how to extract style information directly from style images and fuse it with content images without training by means of an attentional reweighting strategy. Unlike the methods mentioned above, our approach focuses on the interaction of image and text information, as well as a balance for guiding the generation process.
@@
Text-guided style transfer. Stylization [30, 26, 9, 47] is to control the content through text and make the generated image consistent while keeping the style of the reference image. StyleDrop [39] is a method of achieving arbitrary style synthesis from a small number of stylized images and textual descriptions by efficiently fine-tuning a small number of parameters of a pre-trained model and combining iterative training with feedback to improve the quality of the generated images. The well-received work IP-Adapter [48] improves stylization image generation by introducing the embedding of input images in an additional layer of cross-attention, which enhances the model’s ability to capture features from the input images. Building on the IP-Adapter, InstantStyle [42] manually selects specific attention layers to control the style of the output. However, for the IP-Adapter conditioned on natural images, the expected conditioning of the input artistic image does not always work. Visual Style Prompting (VSP) [19] captures stylistic details by fusing key features of the reference image in a later self-attention layer while preserving the content consistency of the original image. However, compared to cross-attention, self-attention provides a weaker ability to control the semantic content of the generated image. Style Aligned [13] attempts to align style and content through a shared attention mechanism. However, it generates results with content information leaked from the style image, and there are challenges in disentangling content and style. StyleShot [10] is trained by a two-stage style control method. However, detailed information within the control is easily lost due to sparse rows and columns. Furthermore, [10], [25], and [46] only trained on the art-text datasets, making it difficult to broadly adapt to arbitrary stylistic features.
3 Method
@@
The overall architecture of ArtCrafter is shown in Fig. 3. As reviewed in Sec. 3.1, ArtCrafter is built upon diffusion model [33, 37]. In Sec. 3.2, we introduce attention-based style extraction, which captures multi-level style information by perceiver attention and multi-layer design. Text-image aligning augmentation (Sec. 3.3) allows the model to dynamically weigh the importance of different parts of the text prompt. This enables a more nuanced and context-aware generation process, resulting in images that are more closely related to the text prompt. Explicit modulation (Sec. 3.4) provides an effective way to combine textual and visual information, enabling the model to generate images that are both relevant to the text prompt and have diverse visual representations.
3.1 Preliminary
@@
Diffusion model [33] consists of two processes: a forward process, which incrementally adds Gaussian noise $\epsilon$ to the data $x_{0}$ through a Markov chain. Additionally, a denoising process generates samples from Gaussian noise $x_{T}\sim N(0,1)$ with a learnable denoising model $\epsilon_{\theta}(x_{t},t,c)$ parameterized by $\theta$. This denoising model $\epsilon_{\theta}(\cdot)$ is implemented with U-Net [34] and trained with a mean-squared loss derived by a simplified variant of the variational bound:
@@
where $c$ denotes an optional condition. In the diffusion model, $c$ is generally represented by the text embeddings $E_{T}$ encoded from a text prompt using CLIP [31] and integrated into the diffusion model through the following designed module (Sec. 3.4).
Figure 4: Qualitative comparison with other state-of-the-art text-guided stylization methods.
3.2 Attention-based Style Extraction
This section elaborates on the style extraction method, which enhances the style encoding capabilities by integrating fine-grained features through a multi-layer architecture. The objective is to capture intricate style details from images by leveraging perceiver attention [18] and position-wise feed-forward network (FFN) [41].
Given the reference image, we obtain the input image embeddings through CLIP, denoted as $x$. The latent variables $z$ are initialized as a tensor with a shape of $(1,N,D)$, where $N$ is the number of queries and $D$ is the dimension of the latent space, and are normalized by dividing by the square root of $D$ to stabilize the training process:
@@ 3.2 Attention-based Style Extraction
The perceiver attention mechanism denoted as P-Attn, is then applied to update the latent variables by attending to the input $x$ and the repeated latent variables:
@@
where $d_{k}$ is the dimension of the key tensor, typically equal to $D$. This operation allows the model to selectively focus on different parts of the input data based on the learnable latent variables.
The FFN consists of two linear transformations with a GELU activation function in between:
$\text{FFN}(z^{\prime})=W_{2}\cdot\text{GELU}(W_{1}\cdot z^{\prime}+b_{1})+b_{2},$ (6)
where $W_{1}$ and $W_{2}$ are weight matrices, and $b_{1}$ and $b_{2}$ are bias terms. The output $E_{I}$ represents the style embeddings extracted from the input image, obtained by combining the FFN output and the updated latent variables $z^{\prime}$:
$E_{I}=\text{FFN}(z^{\prime})+z^{\prime}.$ (7)
3.3 Text-Image Aligning Augmentation
The text-image aligning augmentation method is designed to dynamically prioritize different aspects of the text prompt by leveraging cross-attention mechanisms. This module allows the model to integrate image and text embeddings more effectively, projecting them into a shared feature space where their interactions can be more nuanced.
@@ 3.3 Text-Image Aligning Augmentation
$W_{Q_{\text{I}}}$, $W_{K_{\text{T}}}$, and $W_{V_{\text{T}}}$ are the weight matrices associated with the query for images, key for text, and value for text, respectively. The attention weights $\mathbf{w}_{\text{IT}}$ are calculated by taking the dot product of the query matrix $Q_{\text{I}}$ and the key matrix $K_{\text{T}}$, scaled by the square root of the key dimension $d_{k_{\text{T}}}$ to prevent gradient disappearance:
@@
Using the normalized attention weights $\mathbf{w}_{\text{IT}}$, we compute the weighted sum of the value matrix $V_{\text{T}}$ to generate the multimodal embeddings $E^{\prime}_{\text{IT}}$:
@@
$E^{\prime}_{\text{IT}}$ captures the multimodal context more effectively, allowing the model to generate images that are more closely aligned with the semantic content of the text prompt. This approach is particularly beneficial in scenarios where the textual and visual information needs to be tightly integrated to produce coherent outputs.
3.4 Explicit Modulation
The explicit modulation solves the problem of lack of flexibility in traditional fusion methods by seamlessly fusing the image embeddings with the text-image aligning augmented multimodal embeddings utilizing linear interpolation. Specifically, we fuse the image embeddings $E_{\text{I}}$ with the multimodal embeddings $E^{\prime}_{\text{IT}}$ through linear interpolation:
$E_{\text{F}}=\alpha E_{\text{I}}+(1-\alpha)E^{\prime}_{\text{IT}},$ (12)
where $\alpha$ is a predefined constant controlling the fusion ratio between the original and enhanced embeddings. Ultimately, we concatenate the fused image embeddings $E_{\text{F}}$ with the text prompt embeddings $E_{\text{T}}$ to form the complete prompt embeddings for image generation:
$E_{\text{P}}=E_{\text{T}}\oplus E_{\text{F}},$ (13)
where $\oplus$ denotes the concatenation operation, $E_{\text{P}}$ represents enhanced embeddings and integrates into the diffusion model. By balancing the above embeddings, the model gains a robust and controlled representation that captures multimodal conditions effectively, improving generation performance.
4 Experiments
4.1 Experimental Setup
We train ArtCrafter on a training data consisting of about $500,000$ real image-text pairs from LAION-Aesthetic [36] and $50k$ art-text pairs from our proposed ArtMarket dataset. The images are paired with text descriptions generated by BLIP-2 [22], forming image-text data pairs. More explanations and examples of the data are given in the Supplementary Materials. During both training and inference, we resize the input images to a spatial resolution of $512\times 512$. We have implemented our method over [37] diffusion model. Our training processes are conducted using $8$ NVIDIA A100 GPUs, each with $80GB$ of memory, and a batch size of $8$ per GPU. The inference phase, which consumes $5185$ MiB of memory, takes about one second of sampling time on a single $A100$ at denoising steps of $50$.
@@ 4.1 Experimental Setup
Human-Text $\uparrow$ / 38.89% 35.19% 30.56% 27.78% 24.07% 16.67% 13.89% 11.11% 10.19%
Human-Image $\uparrow$ / 35.19% 15.74% 40.74% 29.63% 25.00% 19.44% 21.30% 12.96% 32.41%
Human-Overall $\uparrow$ / 36.11% 18.52% 37.96% 14.81% 12.04% 22.22% 26.85% 9.26% 33.33%
Table 1: Quantitative comparison. The metrics employed for testing encompass image consistency (including CLIP-Image, DINO-v2, Human-Image), text consistency (CLIP-Text, Human-Text), diversity (LPIPS), and overall quality (Human-Overall) quality.
4.2 Qualitative Evaluations

### v2

@@ I Introduction
Diffusion-based text-to-image generation models [1, 2] have made significant strides in the areas of personalization and customization, particularly in consistent synthesis tasks such as identity protection [3, 4], object customization [5, 6], and style transfer [7, 8, 9, 10, 11, 12, 13]. Among these applications, text-guided style transfer has emerged as a powerful tool, focusing on fine-grained style representation that captures abstract concepts like texture, color, composition, brushstroke, and genre. This approach enables the creation of a diverse range of personalized outputs that are deeply rooted in the semantic essence of the input text.
@@
Current methods for stylization tasks often leverage pre-trained diffusion models [14, 15, 16, 17, 18, 19, 20, 21, 22], enhancing model features by adding a trainable adapter module without full retraining. In text-to-image style transfer applications, adapter-based methods shape the style and content of the output by adjusting the condition guidance scales over the input image and text prompts. However, we have identified three main issues with previous research: 1) Inadequate artistic style representation. Traditional image encoder architectures and training data, which primarily focus on natural images, limit their ability to capture the intricate textures and stylistic nuances of artistic images. 2) Suboptimal text-guided conditions. As shown in Fig. 2, the usual adapter-based method [16] fails to deliver the expected results when using the text condition “Fashion shoes”. This discrepancy arises because the amount of information in image and text embeddings is not equivalent, yet adapter-based methods [14, 15, 16, 17, 18, 19, 20] often directly concatenate the two without addressing the imbalance and disparity, leading to image data overshadowing text prompts during the sampling process. 3) Lack of output diversity. The constrained liberation of textual guidance results in generated outputs that closely resemble the reference images, thereby limiting the diversity of the results.
@@
To address the aforementioned challenges, we introduce ArtCrafter, a novel embedding reframing solution based on diffusion models, specifically tailored for text-guided stylization tasks. ArtCrafter comprises three key components: 1) Attention-based Style Extraction (Sec. III-B): This component leverages multi-level features to capture intricate stylistic details, ensuring more coherent and accurate stylistic encoding. It employs a non-layer refinement module and a multi-layer attention architecture to capture both local and global stylistic elements. Additionally, we introduce the ArtMarket dataset, which pairs art images with descriptive texts, enabling us to fine-tune an encoder initially trained on natural images. This approach retains strong generalization capabilities while being sensitive to the unique visual elements of artistic styles. 2) Text-Image Aligning Augmentation (Sec. III-C): This module enhances the alignment of image and text embeddings through crafted attentional interactions, mapping them into a shared feature space. This alignment ensures that the generated images reflect both the style of the reference image and the content of the textual conditions, thereby enhancing controllability. 3) Explicit Modulation (Sec. III-D): This component enhances the adaptability of conventional fusion techniques through the implementation of linear interpolation and concatenation schemes. This method facilitates the merging of original image and text embeddings with multimodal embeddings, yielding enhanced embeddings. It allows ArtCrafter to generate images that are relevant to the text prompts and exhibit diverse visual representations. Our main contributions are summarized as follows:
@@
  • We introduce ArtCrafter, a lightweight adapter designed to enhance the capabilities of pre-trained diffusion models in text-guided stylized image generation. Our approach focuses on attention-based style feature extraction, effectively capturing both local and global features.
@@
  • We propose an innovative text-image aligning augmentation module that enables robust interaction between reference images and textual descriptions within a shared feature space, significantly enhancing the influence of text prompts on the generative process.
@@
  • The explicit modulation within ArtCrafter optimizes the utilization of multimodal embeddings, offering greater flexibility and diversity than conventional methods. Additionally, ArtCrafter is compatible with additional control conditions and achieves superior performance across various experimental benchmarks compared to state-of-the-art approaches.
@@
II Related Work
II-A Attention Control in diffusion models.
@@ II Related Work
Following the remarkable progress made in the field of pre-training text-to-image diffusion models [2, 1, 23, 24], a series of image editing efforts [25, 26, 27, 28] have emerged. Hertz et al. [29] propose the Prompt-to-Prompt method, which achieves text-based partial image editing and generates edited images that conform to textual conditions by replacing original vocabulary and cross-attention maps. Plug-and-play [30] utilizes the spatial features and self-attention mapping of the original image to guide the diffusion model for text-guided image-to-image translation while preserving the spatial layout of the original image. Later, MasaCtrl [31] proposes a mutual self-attention control technique for coherent image editing. Consistent image editing is achieved by preserving the key and value of the self-attention layer of the source image while conditioning the model with desired text prompts. Recently, StyleID [7] proposes a style migration method without training by injecting styles in the self-attention layer and introduces query preservation, attention temperature scaling, and initial latent AdaIN techniques to minimize the impact of style injection on the original content. $\mathcal{Z}^{*}$ [10] shows how to extract style information directly from style images and fuse it with content images without training by means of an attentional reweighting strategy. Unlike the methods mentioned above, our approach focuses on the interaction of image and text information, as well as a balance for guiding the generation process.
II-B Text-guided style transfer.
@@
Stylization [32, 33, 34, 35] is to control the content through text and make the generated image consistent while keeping the style of the reference image. StyleDrop [36] is a method of achieving arbitrary style synthesis from a small number of stylized images and textual descriptions by efficiently fine-tuning a small number of parameters of a pre-trained model and combining iterative training with feedback to improve the quality of the generated images. The well-received work IP-Adapter [16] improves stylization image generation by introducing the embedding of input images in an additional layer of cross-attention, which enhances the model’s ability to capture features from the input images. Building on the IP-Adapter, InstantStyle [18] manually selects specific attention layers to control the style of the output. However, for the IP-Adapter conditioned on natural images, the expected conditioning of the input artistic image does not always work. Visual Style Prompting (VSP) [37] captures stylistic details by fusing key features of the reference image in a later self-attention layer while preserving the content consistency of the original image. However, compared to cross-attention, self-attention provides a weaker ability to control the semantic content of the generated image. Style Aligned [38] attempts to align style and content through a shared attention mechanism. However, it generates results with content information leaked from the style image, and there are challenges in disentangling content and style. StyleShot [17] is trained by a two-stage style control method. However, detailed information within the control is easily lost due to sparse rows and columns. Furthermore, [17], [14], and [27] only trained on the art-text datasets, making it difficult to broadly adapt to arbitrary stylistic features.
III Method
@@
The overall architecture of ArtCrafter is shown in Fig. 3. As reviewed in Sec. III-A, ArtCrafter is built upon diffusion model [1, 39]. In Section III-B, we introduce attention-based style extraction, which captures multi-level style information by non-layer refinement module and multi-layer design. Text-image aligning augmentation (Section III-C) allows the model to dynamically weigh the importance of different parts of the text prompt. This enables a more nuanced and context-aware generation process, resulting in images that are more closely related to the text prompt. Explicit modulation (Section III-D) effectively combines textual and visual information, enabling the model to generate images that are relevant to the text prompt and have diverse visual representations. In Section III-E, we provide a detailed description of the training and inference processes. During training, we optimize the adapter while keeping the parameters of the pre-trained diffusion model frozen. The adapter is trained on the ArtMarket dataset. During inference, we enhance controllability through classifier-free guidance with dual conditioning, which enables the model to generate images more aligned with the text prompt while retaining the stylistic characteristics of the art image.
III-A Preliminary
The forward diffusion process incrementally adds Gaussian noise $\epsilon$ to the data $x_{0}$ through a Markov chain. Specifically, the data $x_{0}$ is gradually corrupted by noise over $T$ timesteps, resulting in a sequence of progressively noisier data points $x_{0},x_{1},\ldots,x_{T}$. Each step $t$ in the Markov chain is defined by:
$x_{t}=\sqrt{1-\beta_{t}}x_{t-1}+\sqrt{\beta_{t}}\epsilon,$ (1)
where $\beta_{t}$ is a predefined noise schedule that controls the amount of noise added at each step, and $\epsilon\sim\mathcal{N}(0,1)$ is Gaussian noise. By the final step $T$, the data $x_{T}$ is typically indistinguishable from pure Gaussian noise.
The reverse denoising process aims to reconstruct the original data from the noisy data $x_{T}$. This process is driven by a learnable denoising model $\epsilon_{\theta}(x_{t},t,c)$ parameterized by $\theta$. The denoising model is implemented using a U-Net architecture [40], which is capable of capturing complex patterns and dependencies in the data. The denoising process starts from $x_{T}\sim\mathcal{N}(0,1)$ and iteratively refines the data by predicting and removing the noise at each timestep $t$:
$\hat{x}_{t-1}=\sqrt{1-\beta_{t}}\hat{x}_{t}-\sqrt{\beta_{t}}\epsilon_{\theta}(\hat{x}_{t},t,c).$ (2)
The denoising model $\epsilon_{\theta}(\cdot)$ is trained to predict the noise $\epsilon$ added at each step, allowing the model to progressively reconstruct the original data $x_{0}$.
@@ III-A Preliminary
The denoising model $\epsilon_{\theta}(\cdot)$ is trained with a mean-squared loss derived from a simplified variant of the variational bound:
@@
where $c$ denotes an optional condition. In the diffusion model, $c$ is generally represented by the text embeddings $E_{T}$ encoded from a text prompt using CLIP [41]. These text embeddings are integrated into the diffusion model, allowing the model to generate samples conditioned on the provided text prompt. This conditioning mechanism enables the diffusion model to produce outputs that are semantically aligned with the text description, making it a powerful tool for text-to-image synthesis and other conditional generation tasks.
Figure 4: Qualitative Comparison with Other State-of-the-Art Text-Guided Stylization Methods. We conducted a comprehensive qualitative evaluation by comparing our method with various state-of-the-art text-guided stylization methods, including Styleshot [17], Style Aligned [38], VSP [37], InstantStyle [18], InstantStyle (Plus) [18], IP-Adapter [16], IP-Adapter (SDXL) [16], CSGO [27], and StyleCrafter [14].
III-B Attention-based Style Extraction (ASE)
ASE enhances the style encoding capabilities by integrating fine-grained features through a multi-layer architecture. Unlike the methods in [42, 43], which focus on general feature extraction, our ASE method specifically targets intricate style details from images by leveraging Perceiver Attention [42] (P-Attn) and a Non-Linear Refinement Module (NRM). Our technical contributions lie in the novel combination and adaptation of these components to style extraction, which significantly enhances the representation of stylistic characteristics. The method involves several key steps: initializing latent variables, expanding them to match the input batch size, applying P-Attn, and using the NRM to refine the latent variables.
Given the reference image, we obtain the input image embeddings through CLIP, denoted as $x$. The latent variables $z$ are initialized as a tensor with a shape of $(1,N,D)$, where $N$ is the number of queries and $D$ is the dimension of the latent space. To stabilize the training process, the latent variables are normalized by dividing by the square root of $D$:
@@ III-B Attention-based Style Extraction (ASE)
The P-Attn mechanism denoted as P-Attn, is then applied to update the latent variables by attending to the input $x$ and the repeated latent variables:
@@
where $d_{k}$ is the dimension of the key tensor, typically equal to $D$. This operation allows the model to selectively focus on different parts of the input data based on the learnable latent variables, thereby capturing intricate style details from the reference image. Compared to [42, 43], our use of P-Attn in the context of style extraction is novel and specifically tailored to enhance the representation of stylistic nuances.
The Non-Linear Refinement Module (NRM) consists of two linear transformations with a GELU activation function in between:
$\text{NRM}(z^{\prime})=W_{2}\cdot\text{GELU}(W_{1}\cdot z^{\prime}+b_{1})+b_{2},$ (8)
where $W_{1}$ and $W_{2}$ are weight matrices, and $b_{1}$ and $b_{2}$ are bias terms. The NRM further refines the latent variables by applying non-linear transformations, enhancing the representation of the style features. In contrast to [42, 43], our NRM is designed to specifically refine style-related latent variables, resulting in a more detailed and robust style embedding.
The output $E_{I}$ represents the style embeddings extracted from the input image, obtained by combining the NRM output and the updated latent variables $z^{\prime}$:
$E_{I}=\text{NRM}(z^{\prime})+z^{\prime}.$ (9)
This final step integrates the refined latent variables with the non-linear transformations from the NRM, resulting in a robust and detailed style embedding that captures the intricate style details from the reference image. This style embedding can be used to guide the generation process in downstream applications, ensuring that the generated outputs inherit the desired stylistic characteristics.
III-C Text-Image Aligning Augmentation (TIAA)
The TIAA method is designed to dynamically prioritize different aspects of the text prompt by leveraging cross-attention mechanisms. This module allows the model to integrate image and text embeddings more effectively, projecting them into a shared feature space where their interactions can be more nuanced. Our novel process involves three main steps: linear transformation, attention-weight calculation, and attention-weighted value matrix computation.
@@ III-C Text-Image Aligning Augmentation (TIAA)
Here, $W_{Q_{\text{I}}}$, $W_{K_{\text{T}}}$, and $W_{V_{\text{T}}}$ are the weight matrices associated with the query for images, key for text, and value for text, respectively. These linear transformations map the embeddings into a shared feature space, enabling the subsequent attention mechanism to effectively capture the interactions between the image and text representations.
The attention weights $\mathbf{w}_{\text{IT}}$ are calculated by taking the dot product of the query matrix $Q_{\text{I}}$ and the key matrix $K_{\text{T}}$, scaled by the square root of the key dimension $d_{k_{\text{T}}}$ to prevent gradient disappearance:
@@
The normalized attention weights $\mathbf{w^{\prime}}_{\text{IT}}$ represent the importance of each text feature with respect to the image features, allowing the model to dynamically focus on the most relevant parts of the text prompt.
Using the normalized attention weights $\mathbf{w^{\prime}}_{\text{IT}}$, we compute the weighted sum of the value matrix $V_{\text{T}}$ to generate the multimodal embeddings $E^{\prime}_{\text{IT}}$:
@@
The resulting multimodal embeddings $E^{\prime}_{\text{IT}}$ capture the multimodal context more effectively, allowing the model to generate images that are more closely aligned with the semantic content of the text prompt. From our experiments, this approach is particularly beneficial in scenarios where the textual and visual information needs to be tightly integrated to produce coherent outputs. By leveraging cross-attention mechanisms, the model can dynamically prioritize different aspects of the text prompt, leading to more accurate and contextually relevant image generation.
TABLE I: Quantitative Comparison. To provide a comprehensive evaluation of our method against other state-of-the-art text-guided stylization techniques, we conducted extensive quantitative experiments using a diverse set of metrics. These metrics cover key aspects of image quality, including image consistency (measured by CLIP-Image and DINO-v2), text consistency (assessed via CLIP-Text), and diversity (quantified using LPIPS). The best results are highlighted in bold, and the second-best results are marked with underline.
@@
TABLE II: User Study. To further validate the effectiveness of our method from a human perception perspective, we conducted a comprehensive user study. Participants were asked to evaluate the generated images based on three key criteria: content consistency (Human-Content), style consistency (Human-Style), and overall quality (Human-Overall). The best results are in bold while the second-best results are marked with underline.
Human-Content $\uparrow$ 4.27 3.96 3.19 3.85 3.08 2.72 2.92 3.27 2.38 3.42
Human-Style $\uparrow$ 4.08 3.08 3.04 3.58 2.92 3.15 3.81 3.46 2.65 2.81
Human-Overall $\uparrow$ 4.19 3.73 3.23 3.65 2.81 2.96 3.50 3.15 2.54 2.88
III-D Explicit Modulation (EM)
Traditional stylization methods lack flexibility. For example, the adapter-based method in Fig. 2 generates similar results, failing to produce diverse outputs. In contrast, the EM method integrates image embeddings with multimodal embeddings through linear interpolation, flexibly balancing both influences. This significantly enhances the diversity of the output.
Specifically, we fuse the image embeddings $E_{\text{I}}$ with the multimodal embeddings $E^{\prime}_{\text{IT}}$ through linear interpolation:
$E_{\text{F}}=\alpha E_{\text{I}}+(1-\alpha)E^{\prime}_{\text{IT}},$ (14)
where $\alpha$ is a predefined constant controlling the fusion ratio between the original and enhanced embeddings. The value of $\alpha$ can be adjusted to balance the contribution of the image embeddings and the multimodal embeddings, ensuring that the fused embeddings $E_{\text{F}}$ retain the essential features of both modalities.
Ultimately, we concatenate the fused image embeddings $E_{\text{F}}$ with the text prompt embeddings $E_{\text{T}}$ to form the complete prompt embeddings for image generation:
$E_{\text{P}}=E_{\text{T}}\oplus E_{\text{F}},$ (15)
where $\oplus$ denotes the concatenation operation. The resulting prompt embeddings $E_{\text{P}}$ integrate the enhanced multimodal information into the diffusion model. By carefully balancing the contributions of the text and image embeddings, the model gains a robust and controlled representation that effectively captures multimodal conditions, thereby improving the overall generation performance.
III-E Training and Inference
During training, we optimize the adapter while keeping the parameters of the pre-trained diffusion model frozen. The adapter is trained on the ArtMarket dataset, using the same loss function as the original Stable Diffusion (SD):
$\mathcal{L}=\mathbb{E}_{z_{0},\epsilon,c_{t},c_{\text{art}},t}\|\epsilon-\epsilon_{\theta}(z_{t},c_{t},c_{\text{art}},t)\|^{2}.$ (16)
Here, $\epsilon$ is the randomly sampled Gaussian noise, $\epsilon_{\theta}$ is the noise prediction model, and $t$ is the timestep. Note that during training, the latent variable $z$ is constructed with the art image $c_{\text{art}}$ as follows:
$z_{t}=\sqrt{\bar{\alpha}_{t}}\psi(c_{\text{art}})+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (17)
where $\psi(\cdot)$ is the function mapping the original input to the latent space, and $\bar{\alpha}_{t}$ represents the cumulative variance at timestep $t$.
Classifier-free guidance [44] is a well-established technique in the field of diffusion models. During the inference phase, we enhance controllability through classifier-free guidance with dual conditioning. The output at timestep $t$ is given by:
$\hat{\epsilon}_{\theta}(x_{t},c_{t},c_{\text{art}},t)=w\epsilon_{\theta}(x_{t},c_{t},c_{\text{art}},t)+(1-w)\epsilon_{\theta}(x_{t},t),$ (18)
where the classifier-free guidance factor $w$ is set to $0.6$.
IV Experiments
IV-A Experimental Setup
We train ArtCrafter on a training data consisting of about $500,000$ real image-text pairs from LAION-Aesthetic [45] and $50k$ art-text pairs from our proposed ArtMarket dataset. The images are paired with text descriptions generated by BLIP-2 [46], forming image-text data pairs. More explanations and examples of the data are given in the Supplementary Materials. During both training and inference, we resize the input images to a spatial resolution of $512\times 512$. We have implemented our method over the stable diffusion 1.5 version [1]. Our training processes are conducted using $8$ NVIDIA A100 GPUs, each with $80GB$ of memory, and a batch size of $8$ per GPU. The inference phase, which consumes $5185$ MiB of memory, takes about one second of sampling time on a single $A100$ at denoising steps of $50$.
IV-B Qualitative Evaluations